In triangle ABC, AD and BE are medians and G is the centroid. Angle AGE = 30° , AD = 12, BE = 18.  Find area of the triangle

Grade:12th pass

1 Answers

Aditya Gupta
4 years ago
hello sam.
lets find the area of triangle AGE.
we know a centroid divides the median in ratio 2:1. so the length AG= (2/3)*AD= 8. similarly, GE= (1/3)*BE= 6.
now, area of triangle AGE= ½ *AG*GE*sin(angle AGE)= ½ *8*6*sin30= 24*1/2= 12.
now, we know that all medians divide a triangle into six parts of equal area (the proof is quite simple, you can even google this phrase and you will find the proof)
so, the area of the triangle ABC= 6*12= 72
kindly approve :)